# Carbon and Silicon, Coexist or Compete? A Survey on Human-AI Interactions in Agent-based Modeling and Simulation

### v1

@@ Appendix A Appendix
2023 | Zhang et al. (Zhang et al., 2024a) | S-V | Pre-S | Prototype | Agent ![]() |
@@
2024 | SOTOPIA (Zhou et al., 2024b) | None | D-S | Actor | Agent ![]() |
@@
- | PedSim (PedSim, [n. d.]) | S-P | D-S | Observer | Agent ![]() |
@@
- | AnyLogic (Borshchev, 2014) | S-P | D-S | Observer | Agent ![]() |
@@
- | AutoGPT (Significant Gravitas, [n. d.]) | None | D-S | Observer | Agent ![]() |
@@
- | BabyAGI (BabyAGI, [n. d.]) | None | D-S | Observer | Agent ![]() |
@@
2023 | Eloy et al. (Eloy et al., 2023) | S-P | D-S | Actor | Agent ![]() |
@@
2023 | Zubatiy et al. (Zubatiy et al., 2023) | None | D-S | Director | Agent ![]() |
@@
2024 | Jaber et al. (Jaber et al., 2024) | R-P | D-S | Director | Agent ![]() |
@@
- | - | - | Post-S | Observer | Agent ![]() |
@@
2024 | PeerGPT (Liu et al., 2024a) | R-P | D-S | Actor | Agent ![]() |
@@
2024 | ClassMeta (Liu et al., 2024c) | S-P | D-S | Actor | Agent ![]() |
@@
2024 | Attig et al. (Attig et al., 2024) | R-V | D-S | Actor | Agent ![]() |
@@
2024 | Hwang et al. (Hwang and Won, 2024) | R-P | D-S | Director | Agent ![]() |
@@
2024 | DrHouse (Yang et al., 2024) | R-P | Post-S | Observer | Agent ![]() |
@@
2024 | Cuadra et al. (Cuadra et al., 2024) | R-P | D-S | Director | Agent ![]() |
@@
2024 | Sasha (King et al., 2024) | R-P | D-S | Director | Agent ![]() |

### v2

@@ Appendix A Appendix
| 2023 | Zhang et al. (Zhang et al., 2024a) | S-V | Pre-S | Prototype | Agent |
@@
| 2024 | SOTOPIA (Zhou et al., 2024b) | None | D-S | Actor | Agent |
@@
| - | PedSim (PedSim, [n. d.]) | S-P | D-S | Observer | Agent |
@@
| - | AnyLogic (Borshchev, 2014) | S-P | D-S | Observer | Agent |
@@
| - | AutoGPT (Significant Gravitas, [n. d.]) | None | D-S | Observer | Agent |
@@
| - | BabyAGI (BabyAGI, [n. d.]) | None | D-S | Observer | Agent |
@@
| 2023 | Eloy et al. (Eloy et al., 2023) | S-P | D-S | Actor | Agent |
@@
| 2023 | Zubatiy et al. (Zubatiy et al., 2023) | None | D-S | Director | Agent |
@@
| 2024 | Jaber et al. (Jaber et al., 2024) | R-P | D-S | Director | Agent |
@@
| - | - | - | Post-S | Observer | Agent |
@@
| 2024 | PeerGPT (Liu et al., 2024a) | R-P | D-S | Actor | Agent |
@@
| 2024 | ClassMeta (Liu et al., 2024c) | S-P | D-S | Actor | Agent |
@@
| 2024 | Attig et al. (Attig et al., 2024) | R-V | D-S | Actor | Agent |
@@
| 2024 | Hwang et al. (Hwang and Won, 2024) | R-P | D-S | Director | Agent |
@@
| 2024 | DrHouse (Yang et al., 2024) | R-P | Post-S | Observer | Agent |
@@
| 2024 | Cuadra et al. (Cuadra et al., 2024) | R-P | D-S | Director | Agent |
@@
| 2024 | Sasha (King et al., 2024) | R-P | D-S | Director | Agent |